OpenStudy (anonymous):

(1-/x+h)^1(0) - (1/x)^(10)/h

OpenStudy (anonymous):

hmm bad editing i think maybe \[\frac{1}{h}\left(\frac{1}{(x+h)^{10}}-\frac{1}{x^{10}}\right)\]

OpenStudy (anonymous):

if you want the limit as h goes to zero you recoginize this as the derivative of \[\frac{1}{x^{10}}\] which is \[\frac{-10}{x^{11}}\]

OpenStudy (anonymous):

thank you so much for your help. I was stuck on this question for a couple of hours.